OpenStudy (anonymous):

medal Complete the division problem and choose the answer with the correct equation. 14 42 42 = 14 × 3 42 = 14 × 2 42 = 14 × 5 42 = 14 × 4

OpenStudy (alexandervonhumboldt2):

14*x=42 14*x/14=42/14 x=3 thus what do you the answer is